OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (5, 0) and a directrix at x = -5.

OpenStudy (anonymous):

Equation of a parabola with focus \((p,0)\) and directrix \(x=-p\) will have the equation:\[ 4px = y^2 \]

OpenStudy (anonymous):

so, 4(5)(-5)=y^2 @wio

OpenStudy (anonymous):

No, you only need to plug in \(p\), while \((x,y)\) remain as they are.

OpenStudy (anonymous):

okay, so 4(5)x=y^2? 20x+y^2? what now? @wio

OpenStudy (anonymous):

No \(+\). It is \(=\).

OpenStudy (anonymous):

And you're done.

OpenStudy (anonymous):

wow! thats simple! Thank you so much! @wio

OpenStudy (anonymous):

It would be more complex if the vertex wasn't at \((0,0)\).